A car gets 28mpg city and 40mpg highway. It is driven 278 miles and used 8 gallons of gas. How many gallons were used on the highway?
Answer by stanbon(75887)

A car gets 28mpg city and 40mpg highway. It is driven 278 miles and used 8 gallons of gas. How many gallons were used on the highway?
-----------------------
Equation:
miles + miles = 278
28x + 40(8-x) = 278
28x + 320 - 40x = 278
-12x = -42
x = 21/6
x = 7/2 gallons (city driving)
8-(7/2) = 1/2 gallon (highway driving)
=================
Cheers,
Stan H.